A submarine can stand a pressure of

1400 KN/m2


above atmospheric


pressure. Find the maximum depth at

which she can work in sea water of RD

= 1.025.


1
Expert's answer
2021-12-23T16:43:40-0500

By the definition of the pressure, we have:


"P=\rho_{sea\ water}gh=RD_{sea\ water}\rho_wgh,""h=\dfrac{P}{RD_{sea\ water}\rho_wg},""h=\dfrac{1400\times10^3\ \dfrac{N}{m^2}}{1.025\times1000\ \dfrac{kg}{m^3}\times9.8\ \dfrac{m}{s^2}}=139\ m."